OpenStudy (unklerhaukus):

\[(x-2y+1)\text dx+(4x-3y-6)\text dy=0\] \[\frac{\text dy}{\text dx}=\frac{2y-x-1}{4x-3y-6}\]

OpenStudy (unklerhaukus):

\[\frac{\text d X}{ \text d Y}=\frac{2Y+2k-X-h-1}{4X+4h-3Y-3k-6}\]\[[\frac{\text d X}{ \text d Y}=\frac{2Y-X+(2k-h-1)}{4X-3Y+(4h-3k-6)}\]\[(2k-h-1)=0;\qquad(4h-3k-6)=0\]\[\frac{\text d Y}{\text d X}=\frac{2Y-X}{4X-3Y}=\frac{2(Y/X)-1}{4-3(Y/X)}\] \[V=(Y/X);\qquad Y=VX\] \[\frac{\text d Y}{\text d X}=X\frac{\text d V}{\text d X}\] \[X\frac{\text d V}{\text d X}=\frac{2V-1}{4-3V}-V\] \[X\frac{\text d V}{\text d X}=\frac{2V-1}{4-3V}-V=\frac{2V-1}{4-3V}-V\frac{4-3V}{4-3V}\]\[X\frac{\text d V}{\text d X}=\frac{3V^2-2V-1}{4-3V}\] \[\frac{\text dX}{X}=\frac{4-3V}{3V^2-2V-1}\text dV\] \[\ln X=\int \frac{1}{4(V-1)}-\frac{15}{4(3V+1)}\text d V \] \[=\frac{1}{4}\int\frac{1}{V-1}-\frac{15}{3V+1}\text d V\] \[4\ln(X)=\ln(V-1)-5\ln(3V+1)+c\]\[4\ln(X)=\ln(\frac{Y}{X}-1)-5\ln(3\frac{Y}{X}+1)+c\]\[4\ln(x+3)=\ln\left(\frac{y+2}{x+3}-1\right)-5\ln\left(3\frac{y+2}{x+3}+1\right)+c\]\[(x+3)^4={\frac{y+2}{x+3}-1}+\left(3\frac{y+2}{x+3}+1\right)^{-5} +e^c \]

OpenStudy (unklerhaukus):

? did i make a mistake on the last step or somewherelse

OpenStudy (blockcolder):

Pretty sure you have to use product rule in \(\frac{\text{dY}}{\text{dX}}\).

OpenStudy (unklerhaukus):

Thanks, i should have \[\frac{\text d Y}{\text d X}=X\frac{\text d V}{\text d X}+V\] right?

OpenStudy (unklerhaukus):

i haven't written up the partial fractions bit but i think i have got that right

OpenStudy (blockcolder):

Yeah, that's about right.

OpenStudy (unklerhaukus):

the last step where i raised everything the the base e to get rid of the logs, i was a bit confused about what to do with the minus in front of the 5

OpenStudy (blockcolder):

Wait, lemme solve this one.

OpenStudy (blockcolder):

Simplify the 2nd to last line before raising everything to e, then everything will be clear.

OpenStudy (unklerhaukus):

AH! \[4\ln(x+3)=\ln\left((\frac{y+2}{x+3}-1)(3\frac{y+2}{x+3}+1)\right)+c\]

OpenStudy (unklerhaukus):

*\[4\ln(x+3)=\ln\left((\frac{y+2}{x+3}-1)/(3\frac{y+2}{x+3}+1)\right)+c\]

OpenStudy (unklerhaukus):

\[X^4=\frac{V-1}{(3V+1)^5}+c\]

OpenStudy (blockcolder):

You're supposed to multiply by e^c, not add.

OpenStudy (unklerhaukus):

\[4\ln(X)=\ln(V-1)-5\ln(3V+1)+c\]\[4\ln(X)=\ln\left({\frac{(V-1)-5}{(3V+1)^5}}\right)+c\]\[X^4=\frac{V-1}{(3V+1)^5}+e^c\]\[X^4=\frac{(X/Y)-1}{(3(X/Y)+1)^5}+e^c\]\[(x+3)^4=\frac{\frac{x+3}{y+2}-1}{(3\frac{x+3}{y+2}+1)^5}+e^c\]

OpenStudy (blockcolder):

\[e^{\text{blabla}+c}=e^{\text{blabla}}e^c\]

OpenStudy (unklerhaukus):

your saying my last three lines should be \[\dots\times e^c\]

OpenStudy (blockcolder):

Yep. Then replace \(e^c\) with \(A\) to have one constant.

OpenStudy (unklerhaukus):

\[(x+3)^4=\frac{\frac{x+3}{y+2}-1}{(3\frac{x+3}{y+2}+1)^5}\times e^c\] \[(x+3)^4=A\frac{\frac{x+3}{y+2}-1}{(3\frac{x+3}{y+2}+1)^5}\]

OpenStudy (unklerhaukus):

\[\left(3 \frac{x+3}{y+2}+1\right)^5 \times (x+3)^4=A\left( \frac{x+3}{y+2}-1\right)\]

OpenStudy (unklerhaukus):

is there anything else i can do , thank you so much for you help by the way

OpenStudy (blockcolder):

Not really. You can't isolate y+2. Leave it at that.

OpenStudy (unklerhaukus):

Thanks blockcolder cool , now i just have to do 10,11, and 12